OpenStudy (anonymous):

i can help with 10th grade geometry

OpenStudy (anonymous):

ookay ?

OpenStudy (anonymous):

help meeeeeeeeeeeeee pls

OpenStudy (anonymous):

The dockers bridge passes over the coast freeway. The freeway is 20 metres below the bridge and at right angles to it. A car travelling 50 km per hour on the bridge is directly above another car travelling at 100 km per hour on the freeway. Find how fast the cars will be seperating after 1 minute later.

OpenStudy (anonymous):

Solve S = 4v2 for v.

OpenStudy (anonymous):

oh this is not algerbra oh . . .

OpenStudy (anonymous):

@TAMATCHI123 v = 1/2 \[v = 1/2\sqrt{s}\]

OpenStudy (anonymous):

oh so when i do that i will get my answer